OpenStudy (anonymous):

The Sherman family made a down payment of $18, 375 for a new house. They mortgaged the remaining cost of $129,625. The interest paid over 30 years was $273, 116. What was the total cost of the house?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

is 273,116 the total cost paid including principal nd interest?

OpenStudy (anonymous):

yes. thank you!

OpenStudy (anonymous):

downpayment=$18,375 remaining Cost =129,625 Total cost 30 years before was 18375+129625=148,000 After paying interest on the ammount borrowed= 273,116+18375=291,491